### v3

@@ Abstract
This paper presents PipeFusion, an innovative parallel methodology to tackle the high latency issues associated with generating high-resolution images using diffusion transformers (DiTs) models. PipeFusion partitions images into patches and the model layers across multiple GPUs. It employs a patch-level pipeline parallel strategy to orchestrate communication and computation efficiently. By capitalizing on the high similarity between inputs from successive diffusion steps, PipeFusion reuses one-step stale feature maps to provide context for the current pipeline step. This approach notably reduces communication costs compared to existing DiTs inference parallelism, including tensor parallel, sequence parallel and DistriFusion. PipeFusion also exhibits superior memory efficiency, because it can distribute model parameters across multiple devices, making it more suitable for DiTs with large parameter sizes, such as Flux.1. Experimental results demonstrate that PipeFusion achieves state-of-the-art performance on 8L40 PCIe GPUs for Pixart, Stable-Diffusion 3 and Flux.1 models. Our Source code is available at https://github.com/xdit-project/xDiT.
Preliminary work.
@@ 1 Introduction
In recent years, the capabilities of AI-generated content (AIGC) have rapidly advanced, with diffusion models emerging as a leading generative technique in image BlackForestLabs (2024); Esser et al. (2024); Chen et al. (2024a); Peebles & Xie (2023) and video synthesis OpenAI (2024); MetaAI (2024). The model architecture of diffusion models is undergoing a significant transformation. Traditionally dominated by U-Net Ronneberger et al. (2015) architectures, these models are now transitioning to Diffusion Transformers (DiTs) Peebles & Xie (2023). The context length of diffusion models is also increasing, which is crucial for handling long videos and high-resolution images. For example, the current leading open-source image generation model, Flux.1 BlackForestLabs (2024), generates images with a resolution of 1024px (1024x1024), requiring a sequence length of 262 thousand tokens; For 4096px resolution images, the input sequence increases to 4.2 million tokens.
@@
The inference latency of generating long visual sequences from DiTs models is notably high due to the quadratic growth in computation time with the sequence length because of the attention mechanism. Given that a single GPU cannot satisfy the latency requirements for practical applications, it becomes necessary to parallelize the DiTs inference for a single image across multiple computational devices. However, tensor parallelism (TP) Shoeybi et al. (2019) commonly used in serving Large Language Models (LLMs) is inefficient for DiTs due to their large activation volumes. In such cases, the communication costs often outweigh the benefits of parallel computation. To address the challenges posed by long sequences, sequence parallelism (SP) Jacobs et al. (2023); Liu et al. (2023) is applied for the parallel inference of DiTs, which partitions the input image into patches and needs data exchange across different devices in the attention operations.
@@
DiTs inference features input temporal redundancy So et al. (2023); Ma et al. (2024b; a); Yuan et al. (2024); Zhao et al. (2024); MetaAI (2024), indicating a high degree of similarity in both inputs and activations across successive diffusion time steps, which can be leveraged to design more efficient parallel approaches. DistriFusion Li et al. (2024) is a sequence parallelism approach that capitalizes on this characteristic applied for U-Net-based diffusion models. By utilizing local fresh activations in conjunction with stale activations from the previous step, DistriFusion engages in attention and convolution operations. This technique effectively overlaps communication costs within the computation of a diffusion timestep. However, when this method is applied to DiTs, it has a significant drawback of excessive memory usage. DistriFusion maintains the full spatial shape of attention keys (K) and values (V) for all layers. Its memory cost is directly proportional to the sequence length.
We discovered that the potential for DiTs-specific parallelism leveraging input temporal redundancy has not been fully exploited. Existing parallel paradigms, whether TP or SP, require communication of Activations for each DiTs block. In contrast, sequence-level parallel methods like TeraPipe Li et al. (2021), which only need to transfer limited input activations, have not been sufficiently explored. However, TeraPipe is tailored for training LLMs and operates on the assumption that the model uses causal attention. This means that each input token only attends to tokens that come before it. However, DiTs utilize full attention, where each token must attend to all tokens that precede and follow it. Therefore, a direct application of sequence-level pipeline parallelism is not feasible in DiTs. To address this challenge, our paper introduces an innovative approach named PipeFusion that strategically harnesses Input Temporal Redundancy to develop an efficient patch-level pipeline parallelism suitable for DiTs parallel inference. Similar to DistriFusion, when the current pipeline step does not have access to the most recent K and V values, it employs the stable K and V from the previous step, which are approximate to the current ones. This approach allows for more efficient parallel processing in DiTs, capitalizing on the inherent redundancy in input sequences.
@@
PipeFusion demonstrates advantages in communication costs, memory usage, and generation accuracy. Firstly, compared to existing DiTs parallel methods, PipeFusion substantially reduces inter-device communication costs. It only transfers the input of the initial layer and the output of the final layer on each device, whereas other parallel approaches, i.e. sequence and tensor parallel, require to transmit activations for every DiT layer. Secondly, PipeFusion offers lower memory consumption by distributing model parameters across multiple devices. This makes it particularly suitable for DiTs with large model parameters, such as the 12B Flux model BlackForestLabs (2024). In contrast to DistriFusion, PipeFusion reduces the memory buffers required for attention K, V to 1/ ( is the parallel degree). Finally, PipeFusion, by utilizing fresh K and V values over a longer period, achieves higher image generation accuracy compared to DistriFusion, thus being closer to the serial inference results. Experimental results show that PipeFusion achieves the lowest DiTs inference latency on 8L40 PCIe-connected GPUs compared to sequence parallelism, DistriFusion, and tensor parallelism.
Our contributions are as follows:
We systematically investigated the parallel approaches applicable to DiTs, including Tensor Parallel (TP), and Sequence Parallel (SP). We also adapted an asynchronous SP variant DistriFusion from U-Net-based models to DiTs.
We proposed a Patch-level Pipeline Parallel method, PipeFusion, which perfectly combines Input Temporal Redundancy with the TeraPipe. This method achieves the lowest communication cost and reduces the memory consumption of model parameters.
On the latest DiTs—Pixart, Stable Diffusion 3, and Flux.1—PipeFusion demonstrated significant advantages in both inference latency and memory usage.
@@ 2 Background & Related Works
Diffusion Models: Diffusion models utilize a noise-prediction deep neural network (DNN) denoted by to generate a high-quality image. The process starts from pure Gaussian noise and involves numerous iterative denoising steps to produce the final meaningful image , with representing the total number of diffusion time steps. As shown in Figure 1, at each diffusion time step , given the noisy image , the model takes , , uand an additional condition (e.g., text, image) as inputs to predict the corresponding noise within . At each denoising step, the previous image can be obtained from the following equation:
In this context, Update denotes a function that is specific to the sampler, i.e. DDIM Song et al. (2020) and DPM Lu et al. (2022), generally involves operations such as element-wise operations. After multiple steps, we decode from the Latent Space to the Pixel Space using a Variational Autoencoder (VAE) Kingma & Welling (2013). Consequently, the predominant contributor to diffusion model inference latency is attributed to the forward propagation through the model .
Diffusion Transformers (DiTs): The architecture of diffusion model is undergoing a pivotal transition from U-Net Ronneberger et al. (2015) to Diffusion Transformers (DiTs) Jiang et al. (2024); Chen et al. (2023); Ma et al. (2024c), driven by the scaling law demonstrating increased model parameters and training data with enhanced model performance. Unlike U-Nets, which apply convolutional layers capturing spatial hierarchies, DiTs segment the input into latent patches and leverage the transformer’s self-attention mechanism to model relationships within and across these patches.
In DiTs, the input noisy latent representation is decomposed into patches. These patches are embedded into tokens and fed into a series of DiT blocks. DiT blocks generally incorporate Multi-Head Self-Attention, Layer Norm, and Pointwise Feedforward Networks. However, the specifics of different DiT models may vary. For instance, the incorporation of conditioning can be achieved through various methods such as adaptive layer norm Peebles & Xie (2023), cross-attention Chen et al. (2023), and extra input tokens Esser et al. (2024). As diffusion models tackle higher-resolution images and longer visual sequences, they impose a quadratic computational burden on inference.
@@
Input Temporal Redundancy: Diffusion Model entails the iterative prediction of noise from the input image or video. Recent research has highlighted the concept of input temporal redundancy, which refers to the similarity observed in both the inputs and activations across successive diffusion timesteps So et al. (2023); Ma et al. (2024b). A recent study Ma et al. (2024a) further investigates the distribution of this similarity across different layers and timesteps. Based on the redundancy, a branch of research caches the activation values and reuses them in the subsequent diffusion timesteps to prune computation. For example, in the context of the U-Net architecture, DeepCache updates the low-level features while reusing the high-level ones from cache Ma et al. (2024b). Additionally, Tgate caches the cross-attention output once it converges along the diffusion process Ma et al. (2024b). By contrast, in the domain of DiT models, -DiT proposes to cache the rear DiT blocks in the early sampling stages and the front DiT blocks in the later stages Chen et al. (2024b). PAB Zhao et al. (2024) exploits the U-shaped attention pattern to mitigate temporal redundancy through a pyramid-style broadcasting approach. Finally, DiTFastAttn Yuan et al. (2024) identifies three types of redundancies, such as spatial redundancy, temporal redundancy, and conditional redundancy, and presents an attention compression method to speed up generation.
Parallel Inference for Diffusion Model: Given the similar transformer architecture, tensor parallelism (TP) Shoeybi et al. (2019) and sequence parallelism (SP) Jacobs et al. (2023); Liu et al. (2023); Li et al. (2023), which are commonly employed for efficient inference in LLMs, can be adapted for DiTs. TP partitions model parameters across multiple devices, enabling parallel computation and reducing memory requirements on individual devices. However, it necessitates an AllReduce operation for the outputs of both the Attention and Feedforward Network modules, resulting in communication overhead proportional to sequence length. For DiTs with exceptionally long sequences, this communication overhead becomes substantial. In contrast, SP partitions the input image across multiple devices, utilizing methods such as Ulysses or Ring to communicate the Attention input and output tensors. SP demonstrates superior communication efficiency compared to TP, but each device stores the entire model parameters, which can be memory intensive. DistriFusion Li et al. (2024) leverages Input Temporal Redundancy to design an asynchronous sequence parallel method that utilizes stable activations. It is specifically designed for with a U-Net-based models.
3 Methods
@@
In this section, we delve into the application of tensor parallelism (TP), sequence parallelism (SP), and an asynchronous SP variant, namely DistriFusion, for parallel inference in DiTs. Following this, we introduce a novel patch-level pipeline parallel approach that markedly improves communication efficiency compared to existing techniques.
@@
In the remaining of this article, the following notations are used: denotes the sequence length, which corresponds to the number of pixels in the latent space. represents the hidden size of the model. stands for the number of network layers. is the number of computing devices.
3.1 Tensor Parallelism
When the DiT blocks consist of Multi-Head Self-Attention followed by a Feed-Forward Network (FFN), tensor parallelism Shoeybi et al. (2019) (TP) can be applied to this architecture by parallelizing the linear layers in a column-wise manner after a row-wise manner. TP requires two all-reduce operations in both the attention and FFN layers, resulting in a communication overhead of for each layer, as shown in Table 1. This communication overhead is directly proportional to the sequence length, making it particularly significant for long sequences, where the communication cost becomes substantial. For instance, the 0.6B Pixart Model demonstrates poor scalability with Tensor Parallelism (TP), as illustrated in Figure 8. Moreover, for non-standard Transformer structures, such as the Multimodal(MM)-DiT blocks used in Flux and Stable Diffusion 3, arranging column-wise and row-wise weight partitioning presents substantial challenges. Consequently, we did not implement tensor parallelism for these structures.
3.2 Sequence Parallelism
In DiTs, 2D latent images can be flattened into visual token sequences, serving as inputs for transformer computation. Parallel computing along the sequence dimension involves partitioning the image into non-overlapping patches, with each device handling its local patch. The transformer’s attention mechanism Vaswani et al. (2017) necessitates that each token attends to all others. However, each device only has 1/ of the Queries (Q), keys (K), and values (V), necessitating communication to gather intermediate results from other devices.
The current best practices for sequence parallelism are DeepSpeed-Ulysses Jacobs et al. (2023) (SP-Ulysses), Ring Atention Liu et al. (2023) (SP-Ring) or a hybrid of both Fang & Zhao (2024). As shown in Figure 2, SP-Ulysses employs All2All communications to transform the sequence dimension partitioning into partitioning along the hidden size, enabling parallel computation of attention across different heads. SP-Ring is a distributed version of Flash Attention, utilizing peer-to-peer (P2P) transmission of K and V subblocks. The outputs of sequence parallelism are consistent with those obtained from single device computation.
@@ 3.3 DistriFusion: Asynchronous Sequence Parallelism
DistriFusion is an asynchronous sequence parallelism technique that exploits input temporal redundancy, observed as high similarity in inputs and activations across successive diffusion timesteps Li et al. (2024); Ma et al. (2024b). This redundancy allows the use of slightly outdated (stale) activations from the previous timestep instead of relying solely on fresh activations. Despite using stale activations, DistriFusion maintains image generation accuracy with no perceptible loss in quality to the human eye.
While DistriFusion was initially applied to U-Net-based SDXL models Podell et al. (2023), it can be extended to DiT models as well. As illustrated in Figure 2, DistriFusion also operates by partitioning along the sequence dimension. The key distinction, however, lies in its use of asynchronous all-gather operations to collect K and V activations from remote devices. Due to the asynchronous nature of this communication, the activations for the current diffusion step are not immediately available; instead, they become accessible during the next diffusion step.
@@ 3.4 PipeFusion: Patch-level Pipeline Parallel
Existing parallel paradigms, whether Tensor Parallel or Sequence Parallel, require communication of activations for each DiTs block. The sequence-level pipeline parallel method proposed by TeraPipe Li et al. (2021), used in LLMs, can transmit only limited input Activations, and has been proven to be more suitable for long sequence input Qin et al. (2024). However, TeraPipe is designed for transformers using causal attention, where each token only attends to its previous tokens. In contrast, DiTs employ full attention, where each token attends to the computation of both previous and subsequent tokens. Nevertheless, by leveraging Input Temporal Redundancy, we found that TeraPipe can be perfectly applied to DiTs inference. We propose PipeFusion, a patch-level pipelined parallel approach for DiTs, which leverages the input temporal redundancy but exhibits enhanced communication and memory efficiency.
PipeFusion simultaneously partitions the input sequence and the layers of the DiTs backbone as shown at the top of Figure 3. PipeFusion partitions the DiTs model along the direction of data flow. Therefore, each partition contains a set of consecutive layers and is deployed on a GPU. PipeFusion also partitions the input image into non-overlapping patches so that each GPU processes the computation for one patch with its assigned layers in parallel, and the diffusion routine runs efficiently in a pipelined manner. Pipelining requires synchronizations between devices, which would be efficient if the DiT workload is evenly distributed across each GPU. The even partition is easy to achieve as a DiT contains numbers of identical transformer blocks.
Suppose we are currently at diffusion timestep , the previous timestep is as the diffusion process goes along the reversed order of timesteps. The example in Figure 3 demonstrates the pipeline workflow with and , where the activation values of patches at timestep are highlighted. Due to the input temporal redundancy, a device does not need to wait for the receiving of full spatial shape activations for the current timestep to start the computation of its own stage. Instead, it employs the stale activations from the previous timestep to provide context for the current computation. For instance, in micro-step 5, the activations for Patchs P0 and P1 are from timestep , while others are from timestep . In this way, each device maintains the full activations and there is no waiting time after the pipeline is initialized. Bubbles only exist at the beginning of the pipeline. Suppose the total number of diffusion timesteps is , then the effective computation ratio of the pipeline is . As is commonly chosen to be large number for high-quality image generation, the effective computation ratio is high. For example, when and , the ratio is equal to 98.5%.
@@
Similar to DistriFusion, before executing the pipeline, we usually conduct several diffusion iterations synchronously, called warmup steps. During the warmup phase, patches are processed sequentially, resulting in low efficiency. Though the warmup steps cannot be executed in the pipelined manner, the workflow is relatively small compare to the entire diffusion process, and thus, the impact on performance is negligible. We will analyze the impact of warmup and proposed the corresponding solutions in Sec. 6.
PipeFusion simultaneously partitions the image and the model layers to form a pipeline, which is also adopted by a previous method TeraPipe Li et al. (2021). However, there are significant differences between the two methods. Firstly, TeraPipe is designed for training language models and leverages the properties of causal attention, where each token only attends to the attention computation of its previous tokens. In contrast, DiTs employ full attention, where each token attends to the computation of both previous and subsequent tokens. The full attention mechanism requires the usage of input temporal redundancy to construct pipelines. Secondly, TeraPipe adopts complex partition methods over the sequence dimension to achieve load balancing, which is not needed in PipeFusion.

### v4

@@ Abstract
This paper presents PipeFusion, an innovative parallel methodology to tackle the high latency issues associated with generating high-resolution images using diffusion transformers (DiTs) models. PipeFusion partitions images into patches and the model layers across multiple GPUs. It employs a patch-level pipeline parallel strategy to orchestrate communication and computation efficiently. By capitalizing on the high similarity between inputs from successive diffusion steps, PipeFusion reuses one-step stale feature maps to provide context for the current pipeline step. This approach notably reduces communication costs compared to existing DiTs inference parallelism, including tensor parallel, sequence parallel and DistriFusion. PipeFusion enhances memory efficiency through parameter distribution across devices, ideal for large DiTs like Flux.1. Experimental results demonstrate that PipeFusion achieves state-of-the-art performance on 8L40 PCIe GPUs for Pixart, Stable-Diffusion 3, and Flux.1 models. Our source code is available at https://github.com/xdit-project/xDiT.
@@ 1 Introduction
In recent years, the capabilities of AI-generated content (AIGC) have rapidly advanced, with diffusion models emerging as a leading generative technique in image [blackforestlabs2023announcement, esser2024scaling, chen2024pixart, peebles2023scalable] and video synthesis [sora2024, meta2024moviegen]. The model architecture of diffusion models is undergoing a significant transformation. Traditionally dominated by U-Net [ronneberger2015u] architectures, these models are now transitioning to Diffusion Transformers (DiTs) [peebles2023scalable]. The context length of diffusion models is also increasing, which is crucial for handling long videos and high-resolution images.
@@
DiTs suffer from quadratic latency growth in long-sequence generation due to attention computation. Given that a single GPU cannot satisfy the latency requirements for practical applications, it becomes necessary to parallelize the DiTs inference for a single image across multiple computational devices. However, tensor parallelism (TP) [shoeybi2019megatron], commonly used in serving Large Language Models (LLMs), is inefficient for DiTs due to their large activation volumes. In such cases, the communication costs often outweigh the benefits of parallel computation. To address the challenges posed by long sequences, sequence parallelism (SP) [jacobs2023deepspeed, liu2023ring] is applied for the parallel inference of DiTs, which partitions the input image into patches and needs data exchange across different devices in the attention operations.
@@
Recent studies have highlighted the presence of input temporal redundancy in Diffusion Transformers (DiTs) during inference [so2024frdiff, ma2024deepcache, ma2024learning, yuan2024ditfastattn, zhao2024real, meta2024moviegen]. This phenomenon manifests as significant similarity in both inputs and activations across consecutive diffusion time steps, presenting opportunities for optimizing parallel computation approaches. DistriFusion [li2023distrifusion] represents a sequence parallelism method that exploits this characteristic in U-Net-based diffusion models. The approach combines fresh local activations with stale activations from preceding steps during attention and convolution operations, effectively overlapping communication costs within individual diffusion timestep computations. However, applying DistriFusion to DiTs reveals substantial memory inefficiencies, as the method requires maintaining full spatial dimensions of attention keys (K) and values (V) across all layers, resulting in memory costs that scale linearly with sequence length.
@@
To more effectively utilize input temporal redundancy, we propose PipeFusion, a novel patch-level pipeline parallelism. PipeFusion partitions both the DiTs layers and input image patches across GPUs, enabling pipelined parallel processing by leveraging temporal redundancy to reuse stale activations from previous timesteps. Firstly, PipeFusion substantially reduces inter-device communication costs compared to existing DiTs’ parallel methods. It only transfers the input of the initial layer and the output of the final layer on each device, whereas other parallel approaches, i.e. sequence and tensor parallel, require transmitting activations for every DiT layer. Secondly, PipeFusion offers lower memory consumption by distributing model parameters across multiple devices. This makes it particularly suitable for DiTs with large model parameters, such as the 12B Flux.1 model [blackforestlabs2023announcement]. In contrast to DistriFusion, PipeFusion reduces the memory buffers required for attention K, V to 1/ ( is the parallel degree). Finally, PipeFusion, by utilizing fresh K and V values over a longer period, achieves higher image generation accuracy compared to DistriFusion. Experimental results show that PipeFusion achieves the lowest DiTs inference latency on 8L40 PCIe-connected GPUs compared to sequence parallelism, DistriFusion, and tensor parallelism.
@@ 2 Background & Related Works
Diffusion Models and Diffusion Transformers (DiTs): Diffusion models generate high-quality images using a noise-prediction deep neural network (DNN) denoted by . The process begins with pure Gaussian noise and iteratively denoises it through multiple steps to produce the final image , with being the total number of diffusion time steps. As shown in Figure 1, at each step , given the noisy image , the model takes , , and an additional condition (e.g., text, image) as inputs to predict the noise within . The previous image is obtained using:
The function Update is specific to the sampler (e.g., DDIM [song2020denoising], DPM [lu2022dpm]) and involves element-wise operations. Finally, is decoded from the Latent Space to the Pixel Space using a Variational Autoencoder (VAE) [kingma2013auto]. The main contributor to inference latency is the forward propagation through . The architecture of is transitioning from U-Net [ronneberger2015u] to Diffusion Transformers (DiTs) [jiang2024megascale, chen2023pixart, ma2024latte], driven by the scaling law that shows improved performance with increased model parameters and training data. Unlike U-Nets that use convolutional layers to capture spatial hierarchies, DiTs segment the input into latent patches and use the transformer’s self-attention mechanism [vaswani2017attention] to model relationships within and across these patches. The input noisy latent representation is decomposed into patches, embedded into tokens, and fed into a series of DiT blocks, which generally include Multi-Head Self-Attention, Layer Norm, and Feedforward Networks.
@@
Input Temporal Redundancy: Diffusion Model entails the iterative prediction of noise from the input image or video. Recent research has highlighted the concept of input temporal redundancy, which refers to the similarity observed in both the inputs and activations across successive diffusion timesteps [so2024frdiff, ma2024deepcache]. A recent study [ma2024learning] further investigates the distribution of this similarity across different layers and timesteps. Based on the redundancy, a branch of research caches the activation values and reuses them in the subsequent diffusion timesteps to prune computation. For example, in the context of the U-Net architecture, DeepCache updates the low-level features while reusing the high-level ones from cache [ma2024deepcache]. Additionally, Tgate caches the cross-attention output once it converges along the diffusion process [ma2024deepcache]. By contrast, in the domain of DiT models, -DiT proposes to cache the rear DiT blocks in the early sampling stages and the front DiT blocks in the later stages [chen2024delta]. PAB [zhao2024real] exploits the U-shaped attention pattern to mitigate temporal redundancy through a pyramid-style broadcasting approach. Finally, DiTFastAttn [yuan2024ditfastattn] identifies three types of redundancies, such as spatial redundancy, temporal redundancy, and conditional redundancy, and presents an attention compression method to speed up generation.
3 DiT Parallel Methods
@@
In this section, we delve into the application of tensor parallelism (TP), sequence parallelism (SP), and an asynchronous SP leveraging input temporal redundancy, namely DistriFusion, for parallel inference in DiTs. Following this, we introduce a novel patch-level pipeline parallel approach named PipeFusion that markedly improves communication efficiency compared to existing techniques.
@@
In the remainder of this article, the following notations are used: denotes the sequence length, which corresponds to the number of pixels in the latent space. represents the hidden size of the model. stands for the number of network layers. is the number of computing devices.
3.1 Tensor Parallelism & Sequence Parallelism
Tensor parallelism (TP) partitions the linear layers of DiTs column-wise and row-wise. This approach requires two all-reduce operations per layer, incurring a communication overhead of per layer (Table 1). The communication cost scales linearly with the sequence length, leading to poor scalability, especially for large models such as the 0.6B-parameter PixArt model (Figure 8). Additionally, TP poses implementation challenges for non-standard architectures like MM-DiT in Flux/SD3.
Sequence parallelism (SP) partitions images into patches, with each device processing a patch as input. Since attention mechanisms require global interactions, devices must communicate partial query (Q), key (K), and value (V) between each other. As shown in Figure 2, two notable SP implementations are DeepSpeed-Ulysses [jacobs2023deepspeed] (SP-Ulysses) and Ring-Attention [liu2023ring] (SP-Ring). SP-Ulysses uses All2All communication to switch between sequence and hidden-size partitioning during parallel attention, while SP-Ring implements distributed Flash Attention via point-to-point K/V subblock transfers. It is worth noting that SP-Ulysses and SP-Ring can be combined to form USP (Unified Sequence Parallel) [fang2024unified], which is often more efficient than using either method alone.
@@ 3.2 DistriFusion: Asynchronous Sequence Parallelism
DistriFusion is an asynchronous sequence parallelism method that leverages temporal redundancy in the input data. It uses slightly outdated (stale) activations from the previous timestep, rather than requiring fresh activations at each step. Despite this, DistriFusion maintains image generation accuracy without any perceptible loss in quality.
While DistriFusion was initially applied to U-Net-based SDXL models [podell2023sdxl], it can be extended to DiT models as well. As illustrated in Figure 2, DistriFusion also operates by partitioning along the sequence dimension. The key distinction, however, lies in its use of asynchronous all-gather operations to collect K and V activations from remote devices. Due to the asynchronous nature of this communication, the activations for the current diffusion step are not immediately available; instead, they become accessible during the next diffusion step.
@@ 3.3 PipeFusion: Patch-level Pipeline Parallel
PipeFusion simultaneously partitions the input sequence and the layers of the DiTs backbone as shown at the top of Figure 3. PipeFusion partitions the DiTs model along the direction of data flow. Therefore, each partition contains a set of consecutive layers and is deployed on a GPU. PipeFusion also partitions the input image into non-overlapping patches so that each GPU processes the computation for one patch with its assigned layers in parallel, and the diffusion routine runs efficiently in a pipelined manner. Pipelining requires synchronizations between devices, which would be efficient if the DiT workload is evenly distributed across each GPU. The even partition is easy to achieve as a DiT contains number of identical transformer blocks.
Suppose we are currently at diffusion timestep , the previous timestep is as the diffusion process goes along the reversed order of timesteps. The example in Figure 3 demonstrates the pipeline workflow with and , where the activation values of patches at timestep are highlighted. Due to the input temporal redundancy, a device does not need to wait for the receiving of full spatial shape activations for the current timestep to start the computation of its own stage. Instead, it employs the stale activations from the previous timestep to provide context for the current computation. For instance, in micro-step 5, the activations for Patchs P0 and P1 are from timestep , while others are from timestep . In this way, each device maintains the full activations and there is no waiting time after the pipeline is initialized. Bubbles only exist at the beginning of the pipeline. Suppose the total number of diffusion timesteps is , then the effective computation ratio of the pipeline is . As is commonly chosen to be a large number for high-quality image generation, the effective computation ratio is high. For example, when and , the ratio is equal to 98.5%. As long as the patch number is greater than the parallel degree , there will be no bubble generation except for the initial startup overhead. We recommend setting . We have analyzed the impact of on the final performance in Appendix Sec. B.
@@
Similar to DistriFusion, before executing the pipeline, we usually conduct several diffusion iterations synchronously, called warmup steps. During the warmup phase, patches are processed sequentially, resulting in low efficiency. Though the warmup steps cannot be executed in a pipelined manner, the workflow is relatively small compared to the entire diffusion process, and thus, the impact on performance is negligible. We have analyzed the impact of warmup and propose the corresponding solutions in Appendix Sec. B.
PipeFusion theoretically outperforms DistriFusion in terms of generation quality, considering the area of fresh activation. As shown in Figure 4, within a single diffusion timestep, PipeFusion continuously increases the area of fresh activation as the pipeline micro-steps progress from diffusion timestep to . In contrast, throughout the entire diffusion process, DistriFusion constantly maintains one patch of fresh area out of the total patches. Therefore, the results of PipeFusion are closer to the original diffusion method compared to DistriFusion.
The exploitation of input temporal redundancy requires a short warmup phase, since the behavior of diffusion synthesis changes qualitatively over the course of denoising. Accordingly, DistriFusion performs several warmup steps with standard synchronous patch parallelism before switching to its asynchronous mode. For PipeFusion, warmup steps cannot be pipelined and therefore introduce bubbles. However, because the number of warmup steps is typically very small compared to the total diffusion iterations, their overall impact on efficiency is limited. We analyze this effect in Sec. B.
PipeFusion represents a fundamental departure from prior pipeline parallelization methods, such as GPipe [huang2019gpipe] and TeraPipe [li2021terapipe]. Unlike GPipe, which splits along the batch dimension and depends on high concurrency across multiple requests, PipeFusion partitions along the sequence dimension to deliver efficient pipelined execution for a single image. Unlike TeraPipe, which is designed for causal attention in language models, PipeFusion is tailored to the full-attention mechanism of DiTs and exploits temporal redundancy to simultaneously improve efficiency and preserve generation quality.